# Monocular Person Localization under Camera Ego-motion

### v1

@@ Abstract
Localizing a person from a moving monocular camera is critical for Human-Robot Interaction (HRI). To estimate the 3D human position from a 2D image, existing methods either depend on the geometric assumption of a fixed camera or use a position regression model trained on datasets containing little camera ego-motion. These methods are vulnerable to fierce camera ego-motion, resulting in inaccurate person localization. We consider person localization as a part of a pose estimation problem. By representing a human with a four-point model, our method jointly estimates the 2D camera attitude and the person’s 3D location through optimization. Evaluations on both public datasets and real robot experiments demonstrate our method outperforms baselines in person localization accuracy. Our method is further implemented into a person-following system and deployed on an agile quadruped robot.
@@ I INTRODUCTION
Person localization is critical for robotic applications. In practice, accurate and stable person localization not only lays the foundation for robot planning and control in human-robot interaction tasks such as Robot Person Following (RPF) [1], but is also important for egocentric crowd behavior analysis and anomaly detection[2]. In recent years, many methods have been developed for monocular person localization from a moving camera[3, 4, 5], most of which rely on the restrictive assumption that the camera is mounted at a fixed height and attitude (roll and pitch angles) from the ground plane[1]. However, many real-world robotic deployments starkly break this assumption, such as autonomous golf caddies moving on undulating golf courses[6] and quadruped robots traversing rough terrains[7]. Fig. 1 shows our interested scenario where a quadruped robot follows a person through a rugged lawn. In these scenarios, the robot’s fierce ego-motion poses challenges to the stable localization and tracking of the person.
@@
Practically, to localize a person in 3D while avoiding ego-motion perturbations, many robotic applications resort to distance sensors such as UWB[7], LiDAR[8, 9, 10, 11, 12]and RGB-D camera[8]. However, using a monocular camera to realize person localization is worth researching because of its low physical complexity and cost.
@@
Another widely adopted solution against ego-motion is to estimate the robot state[13] to compensate for the camera ego-motion based on extra sensors such as IMU[11, 14] and leg kinematic sensors[9] or use odometry[8, 15] to get the current camera attitude and transform the estimated human location to a stablized robot frame. However, robust and low-drift state estimation is challenging for a highly dynamic robot such as a quadruped, especially for IMU-based odometry, the repetitive impact from a trotting quadruped’s foot contact will bring huge challenge against the odometry drift[16, 17], which introduces accumulated error to person localization.
@@
Therefore, instead of relying on a possibly error-accumulated odometry, one solution is to estimate 3D person location from single-frame observation. This problem is widely researched in the computer vision community[18, 19] and the autonomous driving community[20], including monocular depth estimation[21, 22], human localization[23], 3D human joint estimation[24] and human body mesh recovery[25]. However, most of these methods are based on deep learning, they perform upon the distribution of training data and lack generalizability to scenarios, image projection, or camera ego-motion. Another type of methods represents a human with non-rigid models like SMPL[26] and optimizes the 6-DoF pose of a person’s root joint along with posture and shape parameters. However, their complex non-linear systems with numerous parameters bring challenges to highly real-time applications. We propose to estimate 3D person location from every single image shot. Once per-frame person localization achieves both high precision and real-time efficiency, it enables robust human tracking even under fierce camera ego-motion.
@@
Furthermore, we integrate our method into an RPF system and conduct experiments on a Unitree Go1 quadruped[7] that follows a person through a variety of rough terrains. Our method achieves accurate and stable person localization at a long-term RPF trial with fierce camera ego-motion. Our code and dataset are available at https://medlartea.github.io/rpf-quadruped/.
@@ II RELATED WORK
Estimating a person’s metric depth or 3D location from a monocular image is an ill-posed problem. Existing methods either leverage prior knowledge about the scene, such as planar assumptions or that pedestrians have a known scale, to localize a person through geometric constraints (Sec. II-A), or they employ optimization-based methods (Sec. II-B) or deep-learning-based regression methods (Sec. II-C) to estimate person’s 6-DoF pose. We only consider approaches that can be deployed real-time in robotic systems.
@@ II-A Person Localization with Geometric Model
Existing methods model the human body as single line segment, combined with the assumptions of a ground plane and the upright posture of humans. Based on the correspondence between the detected human joints in the image and the geometric model, these methods solve for the 3D positions.
Fei et al. [27] assumes that all observed people have the same known height and stand upright on the same plane. When at least two individuals are observed, they simultaneously estimate the camera’s pose relative to the horizontal plane and the physical distance between people. Aghaei et al. [28] assumes that the camera is fixed at a known height and a tilt angle within a specific range from the plane. By leveraging the length of the upright human torso, the scale of the model is recovered, enabling the estimation of social distance between people. The aforementioned methods are developed for surveillance perspectives, where the camera is faraway above the plane. Choi et al. [29] proposed a framework for tracking people from a moving monocular camera. He assumes the camera is fixed so the image has a fixed ground horizon line, making it prune to severe pitch and roll changes. Koide et al. [3] proposed the first framework for locating, tracking and following a person using a monocular camera from a robot’s perspective. Ye et al. [4] extended Koide’s method to adapt to partial occlusions by continuously updating an unscented Kalman filter using observable 2D joint information. Bacchin et al. [5] used an omnidirectional camera to extend Koide’s method[3] to equirectangular projection. However, although their methods achieve robust person localization on real-time robot applications, they assume the camera pose relative to the ground plane is fixed, which fails to generalize to platforms with ego-motion and restricts the methods to wheeled robots moving on flat ground.
In summary, existing monocular person localization methods based on geometric model are constrained by the assumption of a fixed camera observing on a flat ground. When the camera is mounted on an agile off-road robot, such as a quadruped, severe ego-motion can violate these geometric assumptions, leading to inaccurate person localization results.
@@ II-B Pose Optimization from Semantic Keypoints
A seminal approach by Pavlakos et al. [30] estimates 6-DoF poses of objects by combining semantic keypoints predicted by CNNs with a deformable shape model. They optimize the 6-DoF pose over the Stiefel manifold using convex relaxation and manifold optimization. While this method achieves robustness across object categories, its reliance on the general rigidity of an object model making it difficult for non-rigid targets like humans. Pavliv et al. [31] applies methods in [30] to track drone swarms using a vision-based headset.
BodySLAM++[32] represents a significant leap by tightly coupling human pose estimation and stereo Visual Inertial Odometry (VIO) in a factor graph. It jointly estimates SMPL body parameters, camera trajectory and inertial states via multi-sensor fusion. Through minimizing error terms dominated by human joint reprojection and also containing shape, posture and motion priors, it optimizes all the parameters by nonlinear least squares in real time. However, while achieving metric-scale accuracy in real time, BodySLAM++ is a multi-image work that depends on extra VIO factors, making deployment costly for resource-constrained robots. Their method’s performance under fierce camera ego-motion is not tested.
@@ II-C Person Localization by Learning-based Regression
Regressing 3D Location from 2D Joints. A prominent line of work leverages 2D human keypoints[23] or semantic patches[33] as intermediate representations to infer 3D positions. MonoLoco[34] and its extension MonoLoco++[23] regress 2D skeletal keypoints to predict probabilistic 3D human locations with uncertainty estimates, addressing the inherent depth ambiguity caused by human height variations. However, these methods heavily rely on benchmark datasets like KITTI[35], which captures street-level perspectives and lack views from a robot with large angle ego-motion, leading to poor generalization in unseen environments. As noted in a recent survey[20], in most monocular 3D detection tasks for autonomous driving scenarios, only the heading angle around the up-axis (yaw angle) is considered, which causes a misalignment problem when the camera suffers ego-motion with non-zero roll/pitch angles.
Learning-based Human Pose Estimation. Another branch of work focuses on end-to-end 3D human pose estimation, reconstructing the spacial location of the body’s root joint along with body posture and body shape from monocular images[18, 19]. However, these methods often prioritize root-relative pose accuracy over absolute pose from the camera. Depth ambiguities persist due to the underconstrained nature of monocular inference, and performance degrades significantly in “in-the-wild” settings lacking precise 3D ground truth[18, 19]. Most of these models require heavy computational resources and adapt to a fixed camera setting, making them impractical for robotic platforms. Advances like Multi-HMR[25] enable singo-shot whole-body mesh and metric depth recovery for multiple humans. It optionally accounts for different camera intrinsics although restricted to ideal pin-hole model. In summary, few works address the joint challenges of absolute localization accuracy, real-time inference and cross-domain generalization.
@@ III METHODOLOGY
We propose an optimization-based person localization method robust to camera ego-motion. By representing a human with a four-point model consisting of neck, hip, knee and ankle, we fit the known 3D human model into the observed 2D image measurements and simultaneously estimate the camera’s 2D attitude (roll and pitch angles) along with the person’s 3D location, which is inspired by the solution for PnP problems based on Bundle Adjustment. We then integrate the proposed localization method into an RPF system, achieving stable person-following behavior on a quadruped robot in rough terrains. We first introduce our human model and 2D observations in Sec. III-A, then we describe our optimization details in Sec. III-B and Sec. III-C. Finally, we introduce our RPF system assisted by the proposed localization method in it in Sec. III-D.
@@ III-A Human Model and Observations
To detect the corresponding 2D projections of on the image plane, firstly, we use YOLOX[36] to detect people’s bounding boxes and AlphaPose[37] to detect 2D human joints within the bounding boxes. We notice AlphaPose has good performance even under occlusion and under image distortion, as evidenced in [4]. The stable 2D joint detection lays the foundation for localization. Secondly, we derive the 2D image coordinates of the four points in by computing the median of the corresponding left and right joints’ image coordinates.
@@ III-B Parameterization and Constraints
Since we model the human body as four collinear points, the pose of the human relative to the camera comprises 2-DoF rotation and 3-DoF translation. Specifically, we ignore the orientation of the human body around its central axis. This simplification makes the problem more tractable while retaining sufficient information for accurate localization.
We assume that a human’s footprint in camera frame represents his position, and the heights of the four points in relative to are known as . In camera frame , the direction of the the human’s central axis is represented by a unit vector from the optical center , as shown in Fig. 2a. Due to the camera’s ego-motion when the robot moves on uneven terrain, is not always parallel to the y-axis of camera frame but has 2-DoF rotation from it, which is defined as the rotation from to , marked as .
@@
where and are elementary rotation matrices of the roll and the pitch angles respectively.
When the person and the robot start moving, the camera rotation and the camera height (the length of segment CR in Fig. 2b) varies due to camera ego-motion, while the person’s location on the virtual plane changes and need to be estimated for velocity control mentioned in Sec. III-D. The parameters of the system comprise a state vector :
@@ III-C Optimization Details
By minimizing , we can optimize variables in . Specifically, we divide into two parts: the translation part and the rotation part . We alternately update one part while fixing the other as Pavlakos[30] did to ease the parameter coupling:
This is a nonlinear least-squares problem with a small number of variables and a dense Jacobian. We can constrain the variables in the state vector with bounds based on the real motion limits of the camera and the person. We employ Dogbox[38], an extension of Powell’s dogleg method[39] to solve this bounded nonlinear least-squares problem. To enhance the robustness of the reprojection error against noise, we employ the Cauchy cost function[40]. The initial values of the state vector are set around a level viewing state at each frame.
@@ III-D Implementation in RPF Framework
Our RPF framework is illustrated in Fig. 3. It is based on our former work[4] and follows a widely adopted RPF pipeline[1, 3, 5, 4] with an image as input and a velocity control command as output. Firstly, we detect human bounding boxes and 2D joints in the image. Secondly, we back-project the measurements into the normalized image plane to make our method generalizable to different cameras. Then we input the normalized measurements into our person localization module, which estimates the person’s footprint location on the virtual plane defined in Fig. 3 by optimization. The localization module contains two phases. In the initialization phase, the robot estimates the target person’s joint heights in a static state (e.g. a quadruped lying on the ground) with known camera attitude and height . Therefore the reprojection error in Eq. 5 becomes:
By minimizing , we can optimize along with the person’s location , which is a linear least squares problem and can be solved by SVD[41]:
In the person-following phase, the robot uses the fixed joint heights in to estimate the person’s location and the camera attitude simultaneously.
After localizing each person, we associate newly estimated locations with the existing tracks in 3D space. Based on a Kalman filter, we update a track with the associated location. We use a person re-ID module[42] to identify the target among all the tracks. Finally, the control module navigates the robot to the person-following position based on the target person’s location .

### v2

@@ Abstract
Localizing a person from a moving monocular camera is critical for Human-Robot Interaction (HRI). To estimate the 3D human position from a 2D image, existing methods either depend on the geometric assumption of a fixed camera or use a position regression model trained on datasets containing little camera ego-motion. These methods are vulnerable to severe camera ego-motion, resulting in inaccurate person localization. We consider person localization as a part of a pose estimation problem. By representing a human with a four-point model, our method jointly estimates the 2D camera attitude and the person’s 3D location through optimization. Evaluations on both public datasets and real robot experiments demonstrate our method outperforms baselines in person localization accuracy. Our method is further implemented into a person-following system and deployed on an agile quadruped robot.
@@ I INTRODUCTION
Person localization is critical for robotic applications. In practice, accurate and stable person localization not only lays the foundation for robot planning and control in human-robot interaction tasks such as Robot Person Following (RPF) [islam2019person], but is also important for egocentric crowd behavior analysis and anomaly detection[anomaly]. In recent years, many methods have been developed for monocular person localization from a moving camera[koide2020monocular, ye2023robot, Bacchin-Italian], most of which rely on the restrictive assumption that the camera is mounted at a fixed height and attitude (roll and pitch angles) from the ground plane[islam2019person]. However, many real-world robotic deployments starkly break this assumption, such as autonomous golf caddies moving on undulating golf courses[caddy] and quadruped robots traversing rough terrains[unitree]. Fig. 1 presents our target scenario: a quadruped robot traversing a rugged lawn while following a person. In such challenging environments, the robot’s severe ego-motion poses challenges to stable person localization and tracking.
@@
Practically, to localize a person in 3D while avoiding ego-motion perturbations, many robotic applications resort to distance sensors such as UWB[unitree], LiDAR[zhang2021efficient, perception_engine, Multi-modal, 3D-pose-korean, leg_tracker], and RGB-D camera[zhang2021efficient]. However, using a monocular camera to realize person localization is worth researching because of its low physical complexity and cost.
@@
Another widely adopted solution against ego-motion is to estimate the robot state[roychoudhury2023perception] to compensate for the camera ego-motion based on IMU[3D-pose-korean, IJSR], leg kinematic sensors[perception_engine], or odometry[zhang2021efficient, aso2021portable] to get the current camera attitude and transform the estimated human location to a stabilized robot frame. However, robust and low-drift state estimation is challenging for a highly dynamic robot such as a quadruped, especially for IMU-based odometry. Specifically, the repetitive foot-ground impacts during a trotting gait induce significant odometry drift[allione2023effects_report, yang2023cerberus], which induces accumulating errors in the person’s estimated location.
@@
Therefore, instead of relying on a possibly error-accumulated odometry, one solution is to estimate 3D person location from single-frame observation. This problem is widely researched in the computer vision community[zheng2023deep, mono-hpe-survey2] and the autonomous driving community[its-survey], including monocular depth estimation[depth_anything, unidepth], human localization[monoloco++], 3D human joint estimation[moon] and human body mesh recovery[multi-hmr2024]. However, most of these methods are based on deep learning; they perform based on the distribution of training data and lack generalizability to scenarios, image projection, or camera ego-motion. Another type of method represents a human with non-rigid models like SMPL[smplify] and optimizes the 6-DoF pose of a person’s root joint along with posture and shape parameters. However, their complex non-linear systems with numerous parameters bring challenges to highly real-time applications. We propose to estimate 3D person location from every single image shot. Once per-frame person localization achieves both high precision and real-time efficiency, it enables robust human tracking even under severe camera ego-motion.
@@
Furthermore, we integrate our method into an RPF system and conduct experiments on a Unitree Go1 quadruped[unitree] that follows a person through a variety of rough terrains. Our method achieves accurate and stable person localization at a long-term RPF trial with severe camera ego-motion. Our code and dataset are available at https://medlartea.github.io/rpf-quadruped/.
@@ II RELATED WORK
Estimating a person’s 3D location from a monocular image is an ill-posed problem due to the inherent ambiguity of 3D-to-2D projection. To address this, existing methods either localize a person using geometric constraints derived from scene priors, such as a planar ground assumption or a known pedestrian scale (Sec. II-A), or directly estimate the person’s full 6-DoF pose via optimization (Sec. II-B) or deep learning-based regression (Sec. II-C). In this work, we focus exclusively on approaches capable of real-time deployment on robotic systems.
@@ II-A Person Localization with Geometric Model
Existing methods model the human body as a single line segment, combined with the assumptions of a ground plane and the upright posture of humans. Based on the correspondence between the detected human joints in the image and the geometric model, these methods solve for the 3D positions.
Fei et al. [fei] assumes that all observed people have the same known height and stand upright on the same plane. When at least two individuals are observed, they simultaneously estimate the camera’s pose relative to the horizontal plane and the physical distance between people. Aghaei et al. [proxemics] assumes that the camera is fixed at a known height and a tilt angle within a specific range from the plane. By leveraging the length of the upright human torso, the scale of the model is recovered, enabling the estimation of social distance between people. The aforementioned methods are developed for surveillance perspectives, where the camera is far away above the plane. Choi et al. [choi2010multiple] proposed a framework for tracking people from a moving monocular camera. He assumes the camera is fixed, so the image has a fixed ground horizon line, making it prone to severe pitch and roll changes. Koide et al. [koide2020monocular] proposed the first framework for locating, tracking, and following a person using a monocular camera from a robot’s perspective. Ye et al. [ye2023robot] extended Koide’s method to adapt to partial occlusions by continuously updating an unscented Kalman filter using observable 2D joint information. Bacchin et al. [Bacchin-Italian] used an omnidirectional camera to extend Koide’s method[koide2020monocular] to equirectangular projection. However, although their methods achieve robust person localization on real-time robot applications, they assume the camera pose relative to the ground plane is fixed, which fails to generalize to platforms with ego-motion and restricts the methods to wheeled robots moving on flat ground.
In summary, existing monocular person localization methods based on geometric models are constrained by the assumption of a fixed camera observing on a flat ground. When the camera is mounted on an agile off-road robot, such as a quadruped, severe ego-motion can violate these geometric assumptions, leading to inaccurate person localization results.
@@ II-B Pose Optimization from Semantic Keypoints
A seminal approach by Pavlakos et al. [6-dof] estimates 6-DoF poses of objects by combining semantic keypoints predicted by CNNs with a deformable shape model. They optimize the 6-DoF pose over the Stiefel manifold using convex relaxation and manifold optimization. While this method achieves robustness across object categories, its reliance on the general rigidity of an object model makes it difficult for non-rigid targets like humans. Pavliv et al. [drone-swarm] applies methods in [6-dof] to track drone swarms using a vision-based headset.
BodySLAM++[henning2023bodyslam++] represents a significant leap by tightly coupling human pose estimation and stereo Visual Inertial Odometry (VIO) in a factor graph. It jointly estimates SMPL body parameters, camera trajectory, and inertial states via multi-sensor fusion. Through minimizing error terms dominated by human joint reprojection and also containing shape, posture, and motion priors, it optimizes all the parameters by nonlinear least squares in real time. However, while achieving metric-scale accuracy in real time, BodySLAM++ is a multi-image work that depends on extra VIO factors, making deployment costly for resource-constrained robots. Their method’s performance under severe camera ego-motion is not tested.
@@ II-C Person Localization by Learning-based Regression
Regressing 3D Location from 2D Joints. A prominent line of work leverages 2D human keypoints[monoloco++] or semantic patches[monois3d] as intermediate representations to infer 3D positions. MonoLoco[monoloco] and its extension MonoLoco++[monoloco++] regress 2D skeletal keypoints to predict probabilistic 3D human locations with uncertainty, addressing the inherent depth ambiguity caused by human height variations. However, these methods heavily rely on benchmark datasets like KITTI[kitti], which captures street-level perspectives and lacks views from a robot with large-angle ego-motion, leading to poor generalization in unseen environments. As noted in a recent survey[its-survey], in most monocular 3D detection tasks for autonomous driving scenarios, only the heading angle around the up-axis (yaw angle) is considered, which causes a misalignment problem when the camera suffers ego-motion with non-zero roll/pitch angles.
Learning-based Human Pose Estimation. These methods aim to reconstruct a human’s posture, shape, and root joint location from a single image[zheng2023deep, mono-hpe-survey2], but they often prioritize root-relative accuracy over the absolute pose relative to the camera. These approaches suffer from depth ambiguities due to the underconstrained nature of monocular inference, and their performance degrades in ’in-the-wild’ settings lacking precise 3D ground-truth [zheng2023deep, mono-hpe-survey2]. Moreover, their high computational cost and dependency on a pre-defined camera model with fixed intrinsics make them impractical for mobile robotic platforms. While recent advances like Multi-HMR [multi-hmr2024] enable multi-person mesh and depth recovery, they are still restricted to ideal pinhole camera models. In summary, few works simultaneously address the challenges of absolute localization accuracy, real-time inference, and cross-domain generalization.
@@ III METHODOLOGY
We propose an optimization-based person localization method robust to camera ego-motion. By representing a human with a four-point model consisting of neck, hip, knee and ankle, we fit the known 3D human model into the observed 2D image measurements and simultaneously estimate the camera’s 2D attitude (roll and pitch angles) along with the person’s 3D location, which is inspired by the solution for PnP problems based on Bundle Adjustment. We then integrate the proposed localization method into an RPF system, achieving stable person-following behavior on a quadruped robot in rough terrain. We first introduce our human model and 2D observations in Sec. III-A, then we describe our optimization details in Sec. III-B and Sec. III-C. Finally, we introduce our RPF system assisted by the proposed localization method in it in Sec. III-D.
@@ III-A Human Model and Observations
To detect the corresponding 2D projections of on the image plane, firstly, we use YOLOX[yolox] to detect people’s bounding boxes and AlphaPose[fang2022alphapose] to detect 2D human joints within the bounding boxes. We notice AlphaPose has good performance even under occlusion and under image distortion, as evidenced in [ye2023robot]. The stable 2D joint detection lays the foundation for localization. Secondly, we derive the 2D image coordinates of the four points in by computing the median of the corresponding left and right joints’ image coordinates.
@@ III-B Parameterization and Constraints
Since we model the human body as four collinear points, the pose of the human relative to the camera comprises 2-DoF rotation and 3-DoF translation. Specifically, we ignore the orientation of the human body around its central axis, which is not critical for localization and can be recovered independently by methods such as Part-HOE [part-hoe]. This simplification makes the problem more tractable while retaining sufficient information for accurate localization.
We assume that a human’s footprint in camera frame represents his position, and the heights of the four points in relative to are known as . In camera frame , the direction of the human’s central axis is represented by a unit vector from the optical center , as shown in Fig. 2a. Due to the camera’s ego-motion when the robot moves on uneven terrain, is not always parallel to the y-axis of camera frame but has 2-DoF rotation from it, which is defined as the rotation from to , marked as .
@@
where and are elementary rotation matrices of the roll and the pitch angles, respectively.
When the person and the robot start moving, the camera rotation and the camera height (the length of segment CR in Fig. 2b) varies due to camera ego-motion, while the person’s location on the virtual plane changes and needs to be estimated for velocity control mentioned in Sec. III-D. The parameters of the system comprise a state vector :
@@ III-C Optimization Details
By minimizing , we can optimize variables in . Specifically, we divide into two parts: the translation part and the rotation part . We alternately update one part while fixing the other as Pavlakos[6-dof] did to ease the parameter coupling:
This is a nonlinear least-squares problem with a small number of variables and a dense Jacobian. We can constrain the variables in the state vector with bounds based on the real motion limits of the camera and the person. We employ Dogbox[dogbox], an extension of Powell’s dogleg method[dogleg] to solve this bounded nonlinear least-squares problem. To enhance the robustness of the reprojection error against noise, we employ the Cauchy cost function[ba]. The initial values of the state vector are set around a level viewing state at each frame.
@@ III-D Implementation in RPF Framework
Our Robot Person Following (RPF) framework is illustrated in Fig. 3. Building upon our previous work[ye2023robot], it follows a standard pipeline [islam2019person, koide2020monocular, Bacchin-Italian, ye2023robot], which takes a monocular image stream as input and produces velocity control commands as output.
The pipeline begins with detecting human bounding boxes and 2D joints from the input image. We back-project the measurements into the normalized image plane to make our method generalizable to different cameras (e.g., pinhole, fisheye, and equirectangular). Subsequently, the normalized measurements are fed into the person localization module, which estimates the person’s footprint location on the virtual plane defined in Fig. 3 by optimization.
The localization module contains two phases. In the initialization phase, the robot estimates the target person’s joint heights in a static state (e.g., a quadruped lying on the ground) with known camera attitude and height . Therefore the reprojection error in Eq. 5 is formulated as:
Minimizing this objective function allows us to jointly optimize for the person’s location and their joint heights . This is a linear least squares problem that can be efficiently solved using SVD[svd]:
In the subsequent person-following phase, the robot utilizes the calibrated joint heights to simultaneously estimate the person’s real-time location and the camera’s attitude (see Sec. III-C).
Following the per-frame localization, a data association step links new detections to existing 3D tracks, each of which is then updated using a Kalman filter. To ensure the robot follows the correct individual, a person re-identification (re-ID) module [re-id] is employed to identify the designated target among all active tracks. Finally, the control module navigates the robot to the person-following position based on the target person’s location .